# A novel 3D food printing technique: achieving tunable porosity and fracture properties via liquid rope coiling

### v1

@@ 3.2.1 Effects of printing height on coiling
In Figure 3a, four series of images display the patterns that emerged from depositing the inks in a circular line with a constant flow rate (9.79 mm3/s), with printing heights ranging from 0.5 cm to 10 cm between the nozzle and platform. To characterize the effect of printing height on coiling properties, we quantified the coiling radius ($R$) and frequency ($\Omega$) for different experiments by analyzing images in Figure 3a.
@@ 3.2.2 Effects of flow rate on coiling
As shown in Figure 4b, the coiling radius does not strongly depend on the flow rate, as it remains relatively constant for all ink types. This is in contrast to the findings of [23], who reported coil radii to decrease as a function of the flow rate, attributed to the shear thinning viscosity of materials upon higher extrusion rates. However, the flow rate range studied here may not have been high enough to investigate similar effects. Figure 4c reveals a linear relationship between the coiling frequency and flow rate, which is congruent with literature [19, 23]. The increasing trend is linear and of a repetitive nature for all samples, but the sample with a lower nozzle diameter ($d=1$ mm) displays a sharper increase. Increasing the coiling frequency via increasing the flow rate (multiple methods to do so are described in SI) is a valuable method for structural designing, as it corresponds to an increase in the number of coils formed per printed distance, and therefore, also the structural porosity.
Figure 4: The effect of flow rate on coiling behavior: (a) sequence of images showing carrot, pea and cookie doughs printed with a 1.5mm nozzle, and cookie dough printed with a nozzle diameter of 1.0mm, both printed at a height of 4 cm and movement speed of 300 mm/min; (b, c) coil radius and frequency as a function of flow rate.
@@ 3.2.3 Optimal printing parameters for coiled designs
For design purposes, it is also noteworthy to mention that the initiation of coiling (moving bed or nozzle) and the printing path (circular versus square) do not affect $\Omega$ and $R$ of the coils (elaborated upon in the SI). Thus, different 3D printers can be used to print coils along different paths while maintaining good reproducibility. However, sharp turns in the printing path, especially at low flow rates, can reverse the coiling direction and should be avoided if undesired.
@@ 3.3.1 Macrostructural porosity
With the identified optimal printing ranges for translated coiling, coiled structures with different degrees of overlap between coils, i.e., porosity, were designed from the cookie and pea inks and were subsequently baked in an oven or air fryer. A selection of structures with increasing macrostructural porosities ($\varnothing_{\text{macro}}$ and $\varnothing_{\text{XRTmacro}}$) and their corresponding X-ray tomographs (XRTs), are showcased in Figure 5a. An overview of all coiled structures and the printer settings used to print them is provided in SI. These images effectively capture the distinctive ‘woven’ nature of the coiled structures, although quite some variation exists between the structures. The low porosity samples appear more tightly woven together, with more overlap among strands that stack upon each other rather uniformly, whereas the higher porosity structures consist of more separated loosely overlapped strands. Yet, pea and cookie structures with similar porosities have similar visual appearances with similar coiling patterns and strand overlap, which can be explained by the similar coiling behavior all inks exhibited (sections 3.2.1 and 3.2.2).
Furthermore, consistent differences were found between $\varnothing_{\text{macro}}$ and $\varnothing_{\text{XRTmacro}}$, indicating that both approaches accurately represent macrostructural porosity (see the SI). However, not all samples were subjected to XRT analyses, so although $\varnothing_{\text{XRTmacro}}$ is preferred, henceforth only $\varnothing_{\text{macro}}$ will be presented. It should also be noted that with the printer settings applied, only a relatively high range of porosities ($0.79\leq\varnothing_{\text{macro}}\leq 0.90$) could be attained. This is a range that is very different from that of conventional cookies or biscuits ($\varnothing_{\text{macro}}\approx 0$) but shows that our LRC-3DFP method allows for the creation of cookie-like products with much higher porosities, potentially affecting the textural and thus sensory properties.
Figure 5: (a) Photographs and reconstructed 3D X-ray tomographs of air fried coiled structures with varying porosity. (b) Top and side-view X-ray tomography scans of post-processed pea and cookie ink strands printed with a 1.5 mm nozzle. The grey areas depict the material matrix and air is represented by the black areas.
@@ 3.3.2 Microstructural porosity and properties
In addition to macrostructural porosity (the air fraction between the filaments), the porosity inside the filaments is defined as microstructural or bulk porosity ($\varnothing_{\text{XRTmicro}}$). From the XRT images in Figure 5b, it can be observed that large visual differences exist between pea and cookie microstructures, although no large differences were found among $\varnothing_{\text{XRTmicro}}$ of the samples, regardless of the printer settings or the ink type (see SI). In pea strands, air is predominantly concentrated in a single large pocket in the center of the filament, whereas air pockets in cookie strands are finer and evenly distributed. All strands were printed with a 1.5 mm nozzle, so from the scale, it can be observed that the pea strand shrunk during post-processing, while the size of the cookie strand remained more or less similar. This can be attributed to the high moisture content of the pea dough (53.10±0.03% wb), resulting in greater water loss during baking, causing it to shrink. The cookie ink, on the other hand, contained less water (15.81±0.15% wb), and the matrix was able to better hold the small gas bubbles, which is known to be a unique property of wheat-based doughs.
@@ 3.4 Textural properties of coiled structures with varying porosities
The post-processed structures were subjected to instrumental cutting and compression tests to measure their mechanical properties. In Figure 6a, the cutting and compression states of a cookie and pea structure are presented. The cross-section of the cuts and the corresponding XRT images are shown in Figures 6b and c for cookie and pea samples, respectively. As the strands of the coiled structures create a network of random-like overlaps, the cross sections show an irregular pattern of individual filament strands, which is also clearly visible in the XRT images (Figure 5).
@@
We will now further discuss the mechanical response of the systems to cutting and compression. To obtain comparable variables among the differently sized samples, the applied force was scaled by the effective area of the sample (including the void) that was in contact with the probe, resulting in pressure (kPa), which is represented as a function of the applied strain (-) in Figures 6d-g. The irregular distribution of strands becomes evident from the response of the system showing a series of individual fracture events that seem to follow a degree of randomness for all samples, indicating that many small fractures occur rather than one large one.
@@
In cuts 1 and 3 in Figure 6d, an increase in pressure during cutting can be observed until an onset of complete fracture is reached at a strain of around 0.4. Figure 6e shows that, upon cutting, structures printed with pea ink present more distinct and jagged peaks and lack a clear point of complete fracture. Upon compression, the cookie samples (Figure 6f) exhibit a continuously increasing fracture pressure regime in low compression (strain of $<0.2$ ), where the pressure does not reduce significantly. However, upon large compression, the pressure falls to almost half of its maximum and fluctuates around that value until the saturation regime is reached for compression strains above  0.5. This is not the case for compression of the pea samples presented in Figure 6g, which present more jagged peaks that fall back to a value of around 0 kPa, indicating the sequential collapsing of individual layers of strands, which can also be observed in Figure 6a.
@@
The differences between the fracture behavior of the cookie and pea systems could be attributed to the fusion of individual filament strands and the differences in the distribution of voids in the filament cross-section. The cookie ink contains fat, causing the individual strands to become more fused during post-processing (Figure 6b), while the individual pea filaments remain separated (Figure 6c) and have thicker cell walls (Figure 5b), leading to more jagged fracture profiles. Furthermore, this irregular fracture behavior of both systems is unlike conventional 3D printed structures or common food samples, where the bulk composition determines the overall fracture behavior of the structure, which is usually prone to complete fracture upon abrupt buckling [36]. This type of singular fracture for solid cookie structures is confirmed in SI.
@@
Even though the individual pressure-strain curves are irreproducible (several repetitions are shown by the color scales), we show that the general character of the fracture events is consistent and primarily depends on the porosity. To further analyze the textural properties of the systems, we quantified their hardness, brittleness, and initial stiffness, as defined in section 2.7, and compared these parameters between different samples. Besides these commonly used parameters in textural analysis, we also quantified the distance between major fractures and the distribution of the meaningful fractures. These analyses aimed to provide a better understanding of the disordered mechanical response and an insight into the sensory perception of coiled structures.
@@
Figure 6: Air fried pea and cookie structures, printed with a nozzle diameter of 1.5 mm, a printing height of 4 cm, and a flow rate of 19.58 mm3/s, resulting in a porosity of 0.81. (a) Stills of both structures being subjected to a cutting test (at  0.5 strain) and a compression test (at  0.25 strain). (b,c) Photographs and XRT scans of cross sections of the cookies (top: cookie, bottom: pea). (d,e) Cutting and compression tests for the cookie sample. (f,g) Cutting and compression tests for the pea sample. The red marks depict the mathematically identified main peaks.
@@ 3.4.1 Hardness
Figure 7: (a) Peak pressure determined from cutting tests as a function of porosity, indicative of hardness. (b) Strain at peak force determined from cutting tests as a function of porosity, giving an indication of brittleness. (c) Initial stiffness of the cookie samples determined from cutting and compression tests, as a function of porosity. (d) $\Delta$strain determined from cutting and compression tests as a function of porosity. All samples were printed with a nozzle with a diameter (d) of 1.0 or 1.5 mm, and were either oven baked (O) or air fried (AF).
@@ 3.4.3 Initial stiffness
The initial stiffness of the cookie samples, represented by the effective Young’s modulus, follows a trend similar to that of the hardness and decreases with increasing structural porosity and decreasing nozzle sizes, as shown in Figure 7c. The initial stiffness obtained from the cutting tests nonlinearly decreases as a function of porosity. However, the initial stiffness based on the compression test remains relatively constant over the entire porosity range. The initial stiffness determined by the compression tests is consistently lower than those obtained from cutting tests due to the different nature of fractures created through cutting and compression. These differences offer insights into anticipating differences in sensory perception of the samples upon biting with frontal teeth or chewing with back teeth, which are expected to display similarities to cutting and compressing, respectively [39]. For the pea samples, no clear trend could be observed for the initial stiffness, although overall, it was higher than that of the cookies. Due to the high standard deviations and resulting irreproducibility, this data was not included in this study, but the graph is presented in SI. Additionally, no significant differences in terms of hardness, brittleness, or initial stiffness were found between nearly all cookie structures baked in an oven and air fryer. This lack of difference in post-processing method can be attributed to the similar macro- and microstructural attributes of the systems.
@@ 3.4.5 Number of fracture events
Crunchiness is usually linked to the number of sensible fracture events, which are characterized by the jaggedness of the fracture profile [40, 41]. In order to understand the crunchiness of the coiled systems upon compression and cutting processes, we investigate the total number of fractures, $N_{\text{tot}}$, for different samples printed with a 1.5 mm nozzle upon cutting and compression. In Figure 8, the total number of fractures as a function of porosity is illustrated. This graph reveals that the pea samples are crunchier than the cookie samples and that compression results in a higher number of fracture events than cutting, which was expected since during the cutting process the stress is localized on a cutting line while during the compression test, it is distributed over a larger region. On average, crunchiness slightly declines with increasing porosity, but considering the deviations in the data that are the result of the random nature of the fractures, this is not generically applicable. Despite the random-like behavior of the system, investigating the fracture force distribution reveals additional details and underlying generic mechanisms.
@@
Figure 8: The total number of fractures as a function of porosity for cookie (blue) and pea (orange) samples upon cutting (dahs) and compression (solid) procedures applied at different sections of the same sample.
3.4.6 Peak distribution
@@
To understand the fracture behavior of coiled structures within a broader context, we analyzed the distribution of the rescaled meaningful fracture force, $d_{f}$. In Figures 9a-d, we illustrate the number of occurrences as a function of their magnitude, divided into 12 bins, for the cookie (compression, Figure 9a; and cutting, Figure 9b) and pea (compression, Figure 9c; and cutting, Figure 9d) samples in a lin-log scale. A generic power-law scaling relationship in the distribution of the fracture force magnitude is given by $\log\left(\frac{N}{N_{tot}}\right)=a+n\cdot d_{f}$, which is known as the Gutenberg–Richter law [42]. This law is a typical characteristic of fracture, and while initially introduced in seismology, it has also been identified in various statistical systems in nature, such as acoustic emission [43, 44]. Moreover, to compare the fracture force distribution across the samples, the slopes of the fitting lines in Figures 9a-d, which represent $n$, are presented as a function of porosity in Figures 9e (compression) and f (cutting).
@@
Comparing the occurrence of fracture magnitudes between cookie and pea samples reveals distinct characteristics that possibly stem from the coiled system’s composition. The cookie samples exhibit a more scattered behavior and a stronger dependence on porosity, particularly evident in the compression experiment (Figure 9a). In contrast, the pea samples demonstrate a similar scaling behavior in both compression and cutting tests (Figures 9c, d). For the cookie samples, the value of $n$ declines slightly as a function of porosity, remaining above -0.05, but significantly decreases to $n\approx-0.2$ in the compression test. However, the scaling relationship exhibits a more consistent behavior for the pea samples, with $n\approx-0.05$ for all samples. These substantial differences may be attributed to the difference in the brittleness of their composition. Our present analysis offers a statistical approach for analyzing fractures in edible systems, which could provide insight into understanding relevant sensory perception.
@@
Figure 9: Frequency distributions of the combined fracture forces of oven-baked and air fried (a, b) cookie and (c, d) pea samples printed with different porosities and a 1.5 mm nozzle. (e, f) The power component in power-low distribution (slope in log-linear plots) for cookie (blue) and pea (orange) samples upon compression (e) and cutting (f) tests. Error bars are obtained based on the deviation of the data from the fitting line. The top row corresponds to the data obtained from compression tests, and the bottom row corresponds to data derived from cutting tests.

### v2

@@ 3.2.1 Effects of printing height on coiling
In Figure 3a, four series of images display the patterns that emerged from depositing the inks in a circular line with a constant flow rate (9.79 mm3/s), with printing heights ranging from 0.5 cm to 10 cm between the nozzle and platform. To characterize the effect of printing height on coiling properties, we quantified the coiling radius ($R$) and frequency ($\Omega$) for different experiments by analyzing images in Figure 3a and associated videos; supplementary videos 1 and 2 exemplify the process of coiling of cookie and pea samples.
@@ 3.2.2 Effects of flow rate on coiling
As shown in Figure 4b, the coiling radius does not strongly depend on the flow rate, as it remains relatively constant for all ink types. This is in contrast to the findings of [25], who reported coil radii to decrease as a function of the flow rate, attributed to the shear thinning viscosity of materials upon higher extrusion rates. However, the flow rate range studied here may not have been high enough to investigate similar effects. Figure 4c reveals a linear relationship between the coiling frequency and flow rate, which is congruent with literature [21, 25]. The increasing trend is linear and of a repetitive nature for all samples, but the sample with a lower nozzle diameter ($d=1$ mm) displays a sharper increase. Increasing the coiling frequency via increasing the flow rate (multiple methods to do so are described in SI C.1) is a valuable method for structural designing, as it corresponds to an increase in the number of coils formed per printed distance, and therefore, also the structural porosity.
Figure 4: The effect of flow rate on coiling behavior: (a) sequence of images showing carrot, pea, and cookie doughs printed with a 1.5 mm nozzle, and cookie dough printed with a nozzle diameter of 1.0 mm, both printed at a height of 4 cm and movement speed of 300 mm/min; (b, c) coil radius and frequency as a function of flow rate.
@@ 3.2.3 Optimal printing parameters for coiled designs
For design purposes, it is also noteworthy to mention that the initiation of coiling (moving bed or nozzle) and the printing path (circular versus square) do not affect $\Omega$ and $R$ of the coils (elaborated upon in SI C.2). Thus, different 3D printers can be used to print coils along different paths while maintaining good reproducibility. However, sharp turns in the printing path, especially at low flow rates, can reverse the coiling direction and should be avoided if undesired.
@@ 3.3.1 Macrostructural porosity
With the identified optimal printing ranges for translated coiling, coiled structures with different degrees of overlap between coils, i.e., porosity, were designed from the cookie and pea inks and were subsequently baked in an oven or air fryer. A selection of structures with increasing macrostructural porosities ($\varnothing_{\text{macro}}$ and $\varnothing_{\text{XRTmacro}}$) and their corresponding X-ray tomographs (XRTs), are showcased in Figure 5a. An overview of all coiled structures and the printer settings used to print them is provided in SI C.3. These images effectively capture the distinctive ‘woven’ nature of the coiled structures, although quite some variation exists between the structures. The low porosity samples appear more tightly woven together, with more overlap among strands that stack upon each other rather uniformly, whereas the higher porosity structures consist of more separated loosely overlapped strands. Yet, pea and cookie structures with similar porosities have similar visual appearances with similar coiling patterns and strand overlap, which can be explained by the similar coiling behavior all inks exhibited (sections 3.2.1 and 3.2.2).
Furthermore, consistent differences were found between $\varnothing_{\text{macro}}$ and $\varnothing_{\text{XRTmacro}}$, indicating that both approaches accurately represent macrostructural porosity (see SI C.4). However, not all samples were subjected to XRT analyses, so although $\varnothing_{\text{XRTmacro}}$ is preferred, henceforth only $\varnothing_{\text{macro}}$ will be presented. It should also be noted that with the printer settings applied, only a relatively high range of porosities ($0.79\leq\varnothing_{\text{macro}}\leq 0.90$) could be attained. This is a range that is very different from that of conventional cookies or biscuits ($\varnothing_{\text{macro}}\approx 0$) but shows that our LRC-3DFP method allows for the creation of cookie-like products with much higher porosities, potentially affecting the textural and thus sensory properties.
Figure 5: Coiled structures: (a) Photographs and reconstructed 3D X-ray tomographs of air fried coiled structures with varying porosity. (b) Top and side-view X-ray tomography scans of post-processed pea and cookie ink strands printed with a 1.5 mm nozzle. The gray areas depict the material matrix, and air is represented by the black areas.
@@ 3.3.2 Microstructural porosity and properties
In addition to macrostructural porosity (the air fraction between the filaments), the porosity inside the filaments is defined as microstructural or bulk porosity ($\varnothing_{\text{XRTmicro}}$). From the XRT images in Figure 5b, it can be observed that large visual differences exist between pea and cookie microstructures, although no large differences were found among $\varnothing_{\text{XRTmicro}}$ of the samples, regardless of the printer settings or the ink type (see SI C.4). In pea strands, air is predominantly concentrated in a single large pocket in the center of the filament, whereas air pockets in cookie strands are finer and evenly distributed. All strands were printed with a 1.5 mm nozzle, so from the scale, it can be observed that the pea strand shrunk during post-processing, while the size of the cookie strand remained more or less similar. This can be attributed to the high moisture content of the pea dough (53.10±0.03% wb), resulting in greater water loss during baking, causing it to shrink. The cookie ink, on the other hand, contained less water (15.81±0.15% wb), and the matrix was able to better hold the small gas bubbles, which is known to be a unique property of wheat-based doughs.
@@ 3.4 Textural properties of coiled structures with varying porosities
The post-processed structures were subjected to instrumental cutting and compression tests to measure their mechanical properties. In Figure 6a, the cutting and compression states of a cookie and pea structure are presented. The cross-section of the cuts and the corresponding XRT images are shown in Figures 6b and c for cookie and pea samples, respectively. As the strands of the coiled structures create a network of random-like overlaps, the cross-sections show an irregular pattern of individual filament strands, which is also clearly visible in the XRT images (Figure 5).
@@
We will now further discuss the mechanical response of the systems to cutting and compression. To obtain comparable variables among the differently sized samples, the applied force was scaled by the effective area of the sample (including the void) that was in contact with the probe, resulting in pressure (kPa), which is represented as a function of the applied strain in Figures 6d-g. The irregular distribution of strands becomes evident from the response of the systems showing a series of individual fracture events that seem to follow a degree of randomness for all samples, indicating that many small fractures occur rather than one large one.
@@
In cuts 1 and 3 in Figure 6d, an increase in pressure during cutting can be observed until an onset of complete fracture is reached at a strain of around 0.4. Figure 6e shows that, upon cutting, structures printed with pea ink present more distinct and jagged peaks and lack a clear point of complete fracture. Upon compression, the cookie samples (Figure 6f) exhibit a continuously increasing fracture pressure regime in low compression (strain of $<0.2$), where the pressure does not reduce significantly. However, upon large compression, the pressure falls to almost half of its maximum and fluctuates around that value until the saturation regime is reached for compression strains above $\approx 0.5$. This is not the case for compression of the pea samples presented in Figure 6g, which present more jagged peaks that fall back to a value of around 0 kPa, indicating the sequential collapsing of individual layers of strands, which can also be observed in Figure 6a.
@@
The differences between the fracture behavior of the cookie and pea systems could be attributed to the fusion of individual filament strands and the differences in the distribution of voids in the filament cross-section. The cookie ink contains fat, causing the individual strands to become more fused during post-processing (Figure 6b), while the individual pea filaments remain separated (Figure 6c) and have thicker cell walls (Figure 5b), leading to more jagged fracture profiles. Furthermore, this irregular fracture behavior of both systems is unlike conventional 3D printed structures or common food samples, where the bulk composition determines the overall fracture behavior of the structure, which is usually prone to complete fracture upon abrupt buckling [38]. This type of singular fracture for solid cookie structures is confirmed in SI D.1.
@@
Even though the pressure-strain curves for all samples display an erratic response and stochastic behavior (with several repetitions indicated by the color scales), we demonstrate that the general characteristics of the fracture events are consistent and primarily depend on their porosity. To further analyze the textural properties of the systems, we quantified their hardness, brittleness, and initial stiffness, as defined in 2.7, and compared these parameters between different samples. Besides these commonly used parameters in textural analysis, we also quantified the distance between major fractures and the distribution of the meaningful fractures. These analyses aimed to provide a better understanding of the disordered mechanical response and an insight into the sensory perception of coiled structures.
@@
Figure 6: Cutting and compression experiments: Air fried pea and cookie structures, printed with a nozzle diameter of 1.5 mm, a printing height of 4 cm, and a flow rate of 19.58 mm3/s, resulting in a porosity of 0.81. (a) Stills of both structures being subjected to a cutting test (at $\approx$0.5 strain) and a compression test (at  0.25 strain). (b,c) Photographs and XRT scans of cross-sections of the cookies (top: cookie, bottom: pea). (d,e) Cutting and compression tests for the cookie sample. (f,g) Cutting and compression tests for the pea sample. The red marks depict the mathematically identified main peaks.
@@ 3.4.1 Hardness
Figure 7: Fracture properties: (a) Peak pressure determined from cutting tests as a function of porosity, indicative of hardness. (b) Strain at peak force determined from cutting tests as a function of porosity, giving an indication of brittleness. (c) Initial stiffness of the cookie samples determined from cutting and compression tests, as a function of porosity. (d) $\Delta$strain determined from cutting and compression tests as a function of porosity. All samples were printed with a nozzle with a diameter (d) of 1.0 or 1.5 mm, and were either oven baked (O) or air fried (AF).
@@ 3.4.3 Initial stiffness
The initial stiffness of the cookie samples, represented by the effective Young’s modulus, follows a trend similar to that of the hardness and decreases with increasing structural porosity and decreasing nozzle sizes, as shown in Figure 7c. The initial stiffness obtained from the cutting tests nonlinearly decreases as a function of porosity. However, the initial stiffness based on the compression test remains relatively constant over the entire porosity range. The initial stiffness determined by the compression tests is consistently lower than those obtained from cutting tests due to the different nature of fractures created through cutting and compression. These differences offer insights into anticipating differences in sensory perception of the samples upon biting with frontal teeth or chewing with back teeth, which are expected to display similarities to cutting and compressing, respectively [41]. For the pea samples, no clear trend could be observed for the initial stiffness, although overall, it was higher than that of the cookies. Due to the high standard deviations and resulting irreproducibility, this data was not included in this study, but the graph is presented in SI D.3. Additionally, no significant differences in terms of hardness, brittleness, or initial stiffness were found between nearly all cookie structures baked in an oven and air fryer. This lack of difference in post-processing method can be attributed to the similar macro- and microstructural attributes of the systems.
@@ 3.4.5 Number of fracture events
Crunchiness is usually linked to the number of sensible fracture events, which are characterized by the jaggedness of the fracture profile [42, 43]. In order to understand the crunchiness of the coiled systems upon compression and cutting processes, we investigate the total number of meaningful fractures, $N_{\text{tot}}$, for different samples printed with a 1.5 mm nozzle upon cutting and compression. In Figure 8, the total number of fractures as a function of porosity is illustrated. This graph reveals that the pea samples are crunchier than the cookie samples. Moreover, compression results in a higher number of fracture events than cutting, which is expected since during the cutting process the stress is localized on a cutting line while during the compression test, it is distributed over a larger region. Overall, crunchiness tends to decline with increasing porosity, but considering the deviations in the data that are the result of the random nature of the fractures, this is not generically applicable. Despite the random-like behavior of the system, investigating the fracture force distribution reveals additional details and underlying generic mechanisms.
@@
Figure 8: The total number of fractures as a function of porosity for cookie (yellow) and pea (green) samples upon compression (solid) and cutting (dash) procedures applied at different sections of the same sample.
3.4.6 Fracture magnitude distribution
@@
To understand the fracture behavior of coiled structures within a broader context, we analyzed the distribution of the rescaled meaningful fracture force, $\Delta f$. In Figures 9a-d, we illustrate the rescaled number of fracture events equal and larger than each fracture magnitude as a function of the magnitude in a linear-log scale, upon compression (cookie, Figure 9a; and pea, Figure 9b) and cutting (cookie, Figure 9c; and pea, Figure 9d) samples. A generic power-law scaling relationship in the distribution of the fracture force magnitude is given by $\log({N}/{N_{\text{tot}}})=a+n\Delta f$, which is known as the Gutenberg–Richter law [44]. This law captures the typical characteristic of fracture, and while initially introduced in seismology, it has also been identified in various statistical systems in nature, such as acoustic emission [45, 46]. Moreover, to compare the fracture force distribution across the samples, the slopes of the fitting lines in Figures 9a-d, which represent $n$, are presented as a function of porosity in Figures 9e (cookie) and f (pea).
@@
Comparing the occurrence of fracture magnitudes between cookie and pea samples reveals distinct characteristics that possibly stem from the coiled system’s composition. The cookie samples exhibit a more scattered behavior and a stronger dependence on porosity, particularly evident in the cutting experiment (Figure 9b). For the cookie samples, the value of $n$ declines as a function of porosity in both compression and cutting tests. However, the pea samples demonstrate a more consistent scaling behavior in both compression and cutting tests (Figures 9b, d), exhibiting $n\approx-0.1$ for all samples. These differences may be attributed to the difference in the brittleness of their composition. The difference in fracture profile can be qualitatively observed in the supplementary videos 3 and 4, respectively, upon compression and cutting tests. Our present analysis offers a statistical approach for analyzing fractures in edible systems, which could provide insight into understanding relevant sensory perception.
@@
Figure 9: Distribution of the fracture magnitude: (a–d) The logarithm of the cumulative occurrence equivalent to or greater than each fracture magnitude, rescaled by the total number of fractures, is illustrated as a function of the dimensionless fracture magnitude for combined oven-baked and air-fried samples during compression (a, cookie; b, pea) and cutting (c, cookie; d, pea) experiments. (e, f) The power component in the power-low distribution (slope in log-linear plots) for cookie (e) and pea (f) samples fractured via compression (solid) and cutting (dash) experiments; error bars are based on the deviation of the data from the fitting line.